If log of x is 2.7756 and log of y is 2.5637, find x+y.

Answer by levi_(1)

log x=2.7756
x=10^2.7756
log y=2.5637
y=10^2.5637
x+y=10^2.7756+10^2.5637
=963(cor to 3sf)